OpenStudy (anonymous):

Simplify: 2sqrt50 - 3sqrt8

OpenStudy (anonymous):

\[4\sqrt{2}\]

OpenStudy (anonymous):

thanks

OpenStudy (anonymous):

\[2\sqrt{50} - 3\sqrt8\] \[=2\sqrt{2\times5^2} - 3\sqrt{2\times2^2}\] \[=2(5)\sqrt{2} - 3(2)\sqrt{2}\] \[=4\sqrt{2}\]

OpenStudy (anonymous):

\[4\sqrt{2}\] will be your answer